Circular No. 8834

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007be–2007bz Numerous supernovae have been discovered on unfiltered CCD frames:

2007be by S. Moretti and S. Tomaselli (Bastia, Ravenna, Italy, 0.42-m reflector); 2007bf and 2007bj by T. Puckett (cf. IAUC 8822; 2007bf with E. Guido, and 2007bj with J. Newton and R. Gorelli); 2007bg by R. Quimby and E. Rykoff (cf. IAUC 8779); 2007bh by T.-w. Chen et al. (cf. IAUC 8667); 2007bi and 2007bo–2007bz by P. E. Nugent and the “Nearby Supernova Factory” collaboration; 2007bk by H. Mikuž (cf. IAUC 8687); and the rest by R. Martin (cf. IAUC 8496). Discovery observations for those objects reported to be brighter than mag 18.7:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007be Apr. 7.96 12^{h}38^{m}06.^{s}65 − 0^{o}01′51′′.4 16.8 18′′.4 E, 14′′.7 S
2007bf Apr. 8.30 14 15 12.05 +15 44 11.8 17.8 4′′.3 E, 20′′.1 S
2007bg Apr. 16.15 11 49 26.18 +51 49 21.8 17.7  —
2007bh Apr. 13.82 15 09 26.37 −10 41 33.2 18.6 24′′.6 E, 10′′.9 N
2007bi Apr. 6.5 13 19 20.19 + 8 55 44.3 18.3  —
2007bj Apr. 18.32 16 22 10.59 − 1 30 51.4 15.1 3′′.9 E, 2′′.3 N
2007bk Apr. 20.00 15 28 45.58 +58 52 12.8 16.5 5′′.0 E, 12′′.5 N
2007bl Apr. 18.86 22 48 05.44 −47 13 54.3 16.5 18′′.9 W, 21′′.6 N
2007bm Apr. 20.68 11 25 02.30 − 9 47 53.8 16.0 2′′.4 W, 10′′.8 S
2007bn Apr. 18.86 22 55 52.81 −63 41 31.6 17.0 27′′ W, 9′′ N
2007bp Mar. 26.4 9 24 00.55 + 5 12 25.0 15.6  —
2007bt Apr. 17.5 14 27 47.73 +12 48 47.1 15.6  —
2007bw Apr. 18.5 17 11 01.99 +24 30 36.4 17.3  —
2007bx Apr. 19.3 11 02 41.52 − 1 38 31.5 18.6  —
2007bz Apr. 22.4 12 56 53.77 +22 22 22.1 16.7  —

Additional information is available on Electronic Telegrams: 2007be in UGC 7800, CBETs 917, 920 (type-II); 2007bf in UGC 9121, CBETs 919, 921 (type-II, discovered a few days past explosion); 2007bg, CBET 927 (type-Ic); 2007bh in PGC 54103, CBET 928; 2007bi, CBET 929 (peculiar type-Ic); 2007bj in NGC 6172, CBETs 930, 932 (type-Ia, discovered perhaps a week past maximum); 2007bk, CBETs 933, 937 (type-Ia); 2007bl in ESO 239-G1, CBET 935; 2007bm in NGC 3672, CBETs 936, 939 (type-Ia, discovered before maximum); 2007bn in NGC 7408, CBET 940; 2007bo– 2007by, CBET 941 (all type-II except type-Ib 2007br); 2007bz in IC 3918, CBETs 941, 943 (type-Ia). Mikuž adds the following R magnitudes for 2007bk: Apr. 21.949 UT, 16.61 ± 0.05; 22.937, 16.76 ± 0.03.

2007 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green